Solve the two equations mvi + MVi = mvf + MVf and vi − Vi = −(vf − Vf) for vf and Vf if m = 1.90 kg, vi = 4.05 m/s, M = 3.10 kg, and Vi = 0.
Answer #1

mvi + MVi = mvf + MVf

1.90 ( 4.05) + 3.10( 0) = 1.90 (vf) + 3.10( Vf)

7.695 = 1.90 (vf) + 3.10( Vf)---------eq(1)

vi − Vi = −(vf − Vf)

4.05 = - vf+ Vf---------eq(2)

vf= Vf- 4.05 ( plugging this value in eq(1)

7.695 = 1.90 (Vf- 4.05) + 3.10( Vf)

7.695 = Vf( 1.90 + 3.10) - 7.695

15.39 = Vf( 5)

Vf= 3.078 m/s

vf= Vf- 4.05 = 3.078 - 4.05= -0.81 m/s